INFRARED SPECTRA OF He-, Ne-, AND Ar-C_{2}D_{2} COMPLEXES

M. REZAI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, AB T2N 1N4, Canada; \underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada; BERTA FERNÁNDEZ, Department of Physical Chemistry and Center for Research in Biological Chemistry and Molecular Materials (CIQUS), University of Santiago de Compostela, E-15782 Santiago de Compostela, Spain; DAVID FARRELLY, Department of Chemistry and Biochemistry, Utah State University, Logan, UT 84322-0300.

Remarkably, there are no previously published experimental spectra of the helium-acetylene van der Waals complex. Apparently, infrared spectra of He-C_{2}H_{2} were recorded around 1990 in Roger Miller’s lab, but a detailed rotational assignment was not possible even with the help of two extensive sets of theoretical predictions.^{a} Here, we study rare gas-C_{2}D_{2} complexes in the ν_{3} region (∼2439 cm^{−1}) using a rapid-scan tuneable diode laser spectrometer to probe a pulsed supersonic slit-jet expansion. The He-C_{2}D_{2} assignment problem is readily apparent: most of the absorption is piled-up in a very narrow region around 2440.85 cm^{−1}, close to the R(0) line of the C_{2}D_{2} monomer. This pile-up is a signature of very weak anisotropy in the helium-acetylene intermolecular potential, leading to almost free internal rotation of the C_{2}D_{2}. We are able to achieve a convincing rotational assignment with the help of theoretical energy level calculations based on the intermolecular potential surface of Munteanu and Fernández.^{b} So far the results are limited to He-C_{2}D_{2} transitions which correlate with the monomer R(0) transition. Ne-C_{2}D_{2} also shows a free-rotation pile-up of lines near R(0) which makes assignment tricky. In contrast, Ar-C_{2}D_{2} exhibits more conventional behavior and a normal asymmetric rotor analysis is possible.

aT. Slee, R.J. Le Roy, and C.E. Chuaqui, Mol. Phys. 77, 111 (1992); R. Moszynski, P.E.S. Wormer, and A. van der Avoird, J.

Chem. Phys. 102, 8385 (1995).

bR. Munteanu and B. Fernández, J. Chem. Phys. 123, 014309 (2005).